OpenStudy (anonymous):

26=6(5-a) How do you get it? Everytime I try I get 29.6

OpenStudy (anonymous):

a=0.666666667

OpenStudy (anonymous):

a = 2/3

OpenStudy (anonymous):

REALLY XDDD

OpenStudy (anonymous):

26/6 = 5-a a = 5 - 26/6 = 2/3